OpenStudy (anonymous):

q=3g+3u. Solve for g. Can someone please show me how to solve this.

OpenStudy (anonymous):

First get all terms with g on one side. Subtract 3u from both sides to do this\[q-3u = 3g\]then divide both sides by 3 to get g by itself\[\frac{q-3u}{3}=g\]

OpenStudy (anonymous):

Thank you.